X+y+z=-7
X+y+2z=3
5x+y+z=-27

Answer by Alan3354(69443)

X+y+z=-7
X+y+2z=3
5x+y+z=-27
-------------
X+y+z=-7
X+y+2z=3
----------- Subtract
-z = -10
z = 10
============
x + y + 10 = -7
x + y = -17
5x+ y = -37 (from 3rd eqn)
-------------- subtract
-4x = 20
x = -5
===============
sub for x & z to find y